A worker's total earnings for one day is $100. He received a $20 fixed payment and consumes 14 hours of leisure. What is the hourly wage rate? Show solutions

1
Expert's answer
2022-02-14T15:33:08-0500

"Total Productive Hours=24-14=10hrs"

"Total Wage=100+20=120" $

"Hourly Wage Rate=\frac{120}{10}=12" $ per hour.